7.The daily delivery of mail at a large city firm follows a time pattern conforming to the Normal

Distribution, with a mean time of arrival of 9.40 am and with a standard deviation of 20 minutes.

Estimate the number of occasions during the 250 working days in the year when the mail arrives

a. before the main gates open at 9 am

b. after the arrival of the office staff at 9.20 am

c. between 10 am and 10.20 am


1
Expert's answer
2021-03-12T06:47:08-0500

Let "X=" difference in time of arrival with respect to 9.40 am in minutes.

"X\sim N(\mu, \sigma^2)." Then "Z=\dfrac{X-\mu}{\sigma}\sim N(0, 1)"

Given "\mu=0, \sigma=20\ min"

a.

"P(X<-40)=P(Z<\dfrac{-40-0}{20})"

"=P(Z<-2)\approx0.02275"

"0.02275(250)=6(occasions)"


b.

"P(X>-20)=1-P(X\leq -20)"

"=1-P(Z\leq\dfrac{-20-0}{20})=1-P(Z\leq-1)""\approx0.84134"

"0.84134(250)=210(occasions)"


c.

"P(20<X<40)=P(X<40)-P(X\leq 20)"

"=P(Z<\dfrac{40-0}{20})-P(Z\leq\dfrac{20-0}{20})"




"=P(Z<2)-P(Z\leq1)"




"\approx0.97725-0.84134\approx0.13591"

"0.13591(250)=34(occasions)"